Assume 5 lb/sec of fluid enter a steady state, steady flow system with p1 = 100 psia, 1 =

0.2 lb/ft3

, v1 = 100 fps, u1 = 800 Btu/lb and leave with p2 = 20 psia, 2 = 0.05 lb/ft3

, v2 = 500

fps, u2 = 780 Btu/lb. During the passage through the open system, 10 Btu/lb of heat is

rejected. Find the work in hp.


1
Expert's answer
2021-11-14T17:12:56-0500

Solution;

From the steady flow equation;

"u_1+\frac{p_1}{\rho_1}+\frac{v_1^2}{2000}+Q=u_2+\frac{p_1}{\rho_1}+\frac{v_2^2}{2000}+W"

Given;

"u_1=800Btu\/lb=1860kJ\/kg"

"P_1=100psia=689.5kPa"

"\rho_1=0.2lb/ft^3=3.2kg/m^3"

"V_1=100fps=30.48m\/s"

"\dot{m}=5lb/s=2.268kg/s"

"u_2=780btu\/lb=1814.18kJ\/kg"

"p_2=20psia=137.895kPa"

"\rho_2=0.05lb/ft^3=0.8kg/m^3"

"V_2=500fps=152.4m\/s"

"Q=10Btu\/kg=23.26kJ\/kg"

Substitute the values into the given equation;

"1860.7+\frac{689.7}{3.2×9.81}+\frac{30.48^2}{2000×9.81}-23.26=W+1814+\frac{137.895}{0.8×9.81}+\frac{152.4^2}{2000×9.81}"

"W=\dot{m}(55.23kJ/kg)"

"W=2.268\u00d755.23=125.27kW"

Convert to horsepower;

168hp